A hydrated compound contains 25.9% calcium, 20.8% sulphur,41.6% oxygen and the rest is water of crystalization. Determine the empirical formula and give its IUPAC name
1
Expert's answer
2019-11-04T07:18:04-0500

Let 100 gm of compound is taken

Then

Moles of Ca taken = "\frac{25.9}{40}=0.6475"

Moles of S taken = "0.65"

Moles of oxygen taken = "\frac{}{}" "\frac{41.6}{16}=2.6"

Moles of water = "\frac{11.7}{18}=0.65"

Dividing all term by 0.65 for getting simple whole no. Ratio

We get,

Ratio of Ca : S : O : "H_2O" = 1 : 1 : 4 : 1

So the compound is

"CaSO_4\u2022H_2O"

IUPAC name = Calcium sulfate